What is the maximum velocity (in m/s) of electrons ejected from a material by 66 nm photons if the work function of the material is 4.43 eV?
1
Expert's answer
2019-09-09T11:08:49-0400

"h" - Plank constant

"\nu" - frequency of photon

"\lambda" - wave lenght of photon

"c" - speed of light

"A" - work function of the material

"m_e" - mass of electron

"V_m" - the maximum velocity of electrons ejected from a material



"\dfrac{c}{\nu}=\lambda \Rightarrow \nu=\dfrac{c}{\lambda}"


Energy of photon "h\nu=\dfrac{hc}{\lambda}" (according to hypothesis of Plank)


According to energy Einstein formula

"h\nu=\dfrac{m_eV_m^2}{2}+A \Rightarrow \dfrac{hc}{\lambda} = \dfrac{m_eV_m^2}{2}+A"


"\dfrac{hc}{\lambda} -A= \dfrac{m_eV_m^2}{2}"


"\dfrac{2hc}{\lambda} -2A= m_eV_m^2"


"\dfrac{2hc}{\lambda} -\dfrac{2A\lambda}{\lambda}= m_eV_m^2"


"\dfrac{2hc-2A\lambda}{\lambda} = m_eV_m^2"


"\dfrac{2(hc-A\lambda)}{\lambda} = m_eV_m^2"


"\dfrac{2(hc-A\lambda)}{\lambda m_e} = V_m^2"


"V_m=\sqrt{\dfrac{2(hc-A\lambda)}{\lambda m_e}}\approx2.25*10^6 \dfrac{m}{s}"


Answer: "2.25*10^6 \dfrac{m}{s}"